1.The length of of a seconds hand in a watch is 1cm.The change in velocity of its tip in 15s is what?


2.A particle is projected vertically upwards and reaches the maxm. ht. H in time T.The ht. of a particle at any time t will be what?

Grade:Upto college level

5 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chandan Das

speed of tip of second hand is constant =2∏*1/60  cm/sec

                                                         =∏/30  cm/sec

so in 15 sec it will rotate by an angel 90 degree

initial velocity = ∏/30 i

final velocity = ∏/30(-j)

change in velocity = | ∏/30(-j)  - ∏/30 i|

                          =∏√2/30  cm/sec


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Nehal Wani
14 years ago
Ans 2: Max Height: u^2sin^?/2g = H usin?/g = T Displacement at any time t = xtan?-(1/2)gx^2/(u^2cos^2?) Replace x with ucos?t : utsin?-(1/2)gt^2 =2Ht/T-(1/2)gt^2
Nehal Wani
14 years ago

Max Height Attained : H = (u2sin2θ)/2g      ½

 

Time Taken For Achieving Max. Height : T = usinθ/g

Displacement At Any time t: xtanθ-½gx2/(u2cos2θ)

Replace x with ucosθt : utsinθ-½gt2

= 2Ht/T-½gt2

 

Meera
6 years ago
What is the answer jjfhhvhkbdriknbvfdtuinnnbbnnhnnnfdchjnnmnnbvfgghjkkkjjfhhvhkbdriknbvfdtuinnnbbnnnkyhjnnnnkkkkkrkknbvfgghjkkk and zoology notes and put
ankit singh
3 years ago
The hand moves 360 degrees in 60 seconds. So in 15 seconds it moves 90 degrees. The tip of the hand is moving around the circumference of the circle with radius 1cm = 2pi cm. It does this in 60 seconds so its speed is: 2 x pi x r / 60 cm per seconds = 0.105 cm per second.